# Universal coarse geometry of spin systems

### v1

@@ Abstract.
The prospect of realizating highly entangled states on quantum processors with fundamentally different hardware geometries raises the question: to what extent does a state of a quantum spin system have an intrinic geometry? In this paper, we propose that both states and dynamics of a spin system have a canonically associated coarse geometry in the sense of Roe on the set of sites in the thermodynamic limit. For a state $\phi$ on an (abstract) spin system with an infinite collection of sites $X$, we define a universal coarse structure $\mathcal{E}_{\phi}$ on the set $X$ with the property that a state has decay of correlations with respect to a coarse structure $\mathcal{E}$ on $X$ if and only if $\mathcal{E}_{\phi}\subseteq\mathcal{E}$. We show that under mild assumptions, the coarsely connected completion $(\mathcal{E}_{\phi})_{con}$ is stable under quasi-local perturbations of the state $\phi$. We also develop in parallel a dynamical coarse structure for arbitrary quantum channels, and prove a similar stability result. We show that several order parameters of the state only depend on the coarse structure of the metric, and we establish a basic compatibility between the dynamical coarse structure associated to a quantum circuit $\alpha$ and the coarse structure of the state $\psi\circ\alpha$ where $\psi$ is any product state.
@@ 1. Introduction
Recently, there have been remarkable experimental realizations of highly entangled, topologically ordered ground states of gapped quantum spin liquids on quantum processors ([S+21, I+24, Xo24, A+23]). The states constructed in these experiments have (at least in principle) the same entanglement structure as the ground state of the theoretical spin liquid up to local errors. Thus, properties of the spin liquid which emerge from the entanglement structure of the ground state alone, and are robust against local perturbation, are arguably realized in the processor rather than merely simulated. A key example of such a property is topological order [Kit06, ZCZW19],
However, most of the interesting properties that we think of as emerging from the entanglement structure of a state also depend on the geometry of the lattice sites to even define mathematically. In the case where the spatial geometry of the processor’s hardware roughly agrees with the geometry of the system being realized, this presents no issues. However, for many states of interest, the geometry underlying the processor’s hardware is neccessarily incompatible with the geometry of the system from which the state originates (which occurs to some extent in the experiment [I+24]).
To illustrate the problem, suppose that we prepare the ground state of a 3D spin liquid (e.g. a Walker-Wang model [WW12]) on a quantum processor with 2D hardware, making use of some (necessarily non-local) bijective encoding scheme to identify sites of the 3D lattice with sites in the 2D processor111in principle we could imagine a more complicated encoding scheme which doesn’t directly map sites to sites, but we will see that allowing this can dramatically change the universal features of a system. If we, as the experimenter, remember this encoding scheme, then it seems that the geometric incompatibility between the two systems may not be an issue, since the universal properties depending on the 2D metric of the system are still mathematically well defined as long as we utilize our encoded 3D metric. However, if we are simply handed the state on the processor and not told the encoding scheme, is it possible in principle to work out the “correct” geometry on the set of sites with which to study and measure properties such as topological order, independently of the physical spatial relations of the qubits?
@@
Is there a precise mathematical sense in which the “correct” geometry on the set of sites of a spin system, with respect to which other entanglement-based order parameters should be defined, is intrinsic to a state? Or do we need to rely on the extrinsic spatial geometry of the sites to define entanglement-based order parameters?
Another theoretical motivation for the above question comes from making sense of various notions of equivalence of quantum phases. Traditionally, we say ground states of two gapped Hamiltonians are in the same quantum phase if there is a (smooth, continuous, etc.) path of Hamiltonians between the two that keeps the gap open. However, the quantum information perspective has motivated the consideration of state-based definitions of equivalence that do not require reference to a parent Hamiltonian. Examples include finite-depth circuit equivalence, locality preserving equivalence [CGW10, ZCZW19], and quasi-local automorphic equivalence [BMNS12]. We stress that all of these notions depend not just on the state, but a-priori also on the underlying geometry of the lattice to define. Thus the“intrinsic geometry” of a state should tell you which geometric structure you should be using when trying to characterize quantum phases.
@@
For a fixed set $X$, it will be useful for us to think of the collection of all coarse structures on it as a partially ordered set with respect to containment. Now note that two metrics on $X$ induce the same coarse structure precisely when they have the same controlled set, and having the same coarse structure will the notion that is most relevant to use, rather than abstract coarse equivalence.
@@
$(\mathcal{E}_{\phi})_{con}$ in the Theorem above can be viewed as a “connected completion” of $\mathcal{E}_{\phi}$ generated by adding all finite subsets of $X\times X$ to $\mathcal{E}_{\phi}$.
The above theorem gives a complete characterization of (connected) coarse structures with respect to which $\phi$ is local, and gives natural candidates for a choice of “universal coarse geometry”. There are other natural variations of $\mathcal{E}_{\phi}$ that are similarly universal but with additional hypotheses. The main variation we consider in this paper involves higher order controlled decay, which results gives rise to the universal coarse structure $\mathcal{E}_{\widetilde{\phi}}$ (see Definition 4.7 and Theorem 4.8). In this case, instead of considering operators localized at points, we consider operators localized in uniformly bounded regions. For resonable nice states, we would expect all these variations to agree. The reason we consider all these variations here is that they make the statement of some of our theorems more convenient.
@@ 1.2. Stability under quasi-local perturbation
Let $A$ be an (abstract) spin system over the set $X$, and $\alpha:A\rightarrow A$ be any ucp map such that $\mathcal{E}_{\alpha}$ is proper and monogenic. Then for any quasi-local perturbation $\Psi:A\rightarrow A$, $(\mathcal{E}_{\phi\circ\Psi})_{con}=(\mathcal{E}_{\phi})_{con}$.
@@ 1.4. Coarse dependence of order parameters
In the final section of the paper, we consider the observation that many entanglement-based order parameters of a state, such as topological order, depend explicitly on some kind of geometry to define. We have proposed (a family of) candidates for canonical coarse geometries that need not be compatible with spatial geometries directly, but we need to work out the extent that order parameters of interest only depend on coarse geometries in the first place. This is the topic of Section 7.
@@
The goal, then, is to which definitions of various order parameters, depending a-priori on a metric $d$, are invariant under this equiavalence relation.
Superselection sectors and topological order. The first order parameter we investigate in Section 7.1 is the W*-category of cone-localized superselection sectors of a state $\phi$ defined on a spin system in $\mathbbm{Z}^{n}$. In 2-dimensions, under some assumptions this category can be upgraded to a braided tensor category, which characterizes the states’ topological order [Naa11, Oga22]. For $(X,d)$ having finite asymptotic dimension (see Definition 4.9), we give a coarse variation of the definition of superselction sectors in terms of coarse cones, and we show this W*-category only depends on the coarse structure. We also show that in the case that the space is $\mathbbm{Z}^{n}$ with the Euclidean metric, this agrees with the usual definition of superselction sectors in terms of localization in honest cones (see Proposition 7.4), hence this W*-category is a coarse order parameter.
@@
(c.f. Proposition 7.6). The definition correlation length 00, finite correlation length, and infinite correlation for a state $\phi$ each depend only on the quasi-equivalence class of the metric. Thus if $\mathcal{E}_{\phi}$ is monogenic, then having 00, finite, or infinite correlation length is a coarse order parameter.
@@ 1.7. Acknowledgements
We would like to thank Dominic Williamson and David Penneys for interesting and useful discussions around these topics. Both authors were supported by NSF DMS- 2247202.
@@ 2. C*-algebra formalism
The mathematical setup for this paper uses the C*-algebra framework for describing physical systems. We briefly review this here, but refer the reader to [BR87, Lan17] for all definitions and results we mention below concerning C*-algebras and a more extensive overview of their role in mathematical physics.
@@
A single automorphism $\alpha\in Aut(A)$ is interpreted as a discrete time evolution. If a system is in state $\phi$ at time 00, then after one time step, it is in state $\phi\circ\alpha$. A continuous time-evolution is decribed by a homomorphism $\mathbbm{R}\rightarrow\text{Aut}(A)$, $t\mapsto\alpha_{t}$. In this case, if a system is initially in state $\phi$ at time $t=0$, then at time $t$ the state is $\phi\circ\alpha_{t}$. If the homomorphism is continuous in the point norm topology, $\{\alpha_{t}\}$ has an infinitesimal generator $\delta$ which is a (generallly unbounded) derivation on the C*-algebra $A$. This can be identified as the Hamiltonian of the system, and the equation expressing that time evolution is generated by $\delta$ is a generalization of the Schrodinger equation (for further details see [BR87]).
More generally, we may consider open dynamics, which are given instead by quantum channels. These are expressed in the Heisenberg picture as unital, completely positive maps $\alpha:A\rightarrow A$, and the action on states is $\phi\mapsto\phi\circ\alpha$ (see [Pau02]). Clearly, automorphisms (which we will call “unitary dynamics”) give examples of quantum channels. However, channels actually simultaneously generalize states and dynamics, since any state also defines a channel, via the map $a\mapsto\phi(a)1_{A}$. In this paper, we will be thinking of a unital completely positive maps mostly as a generalized dynamics.
@@ 2.1. Discrete nets of C*-algebras
. The fundamental example of a quantum system where the algebraic framework can be very useful is the thermodynamic limit of a spin system [BR97]. Here, we start with a (usually countably infinite) set of sites $X$. At each site $x\in X$ we have a unital C*-algebra $A_{x}$ of localized observables. In this paper, we will assume this algebra is nuclear, so that we can ignore the different types of tensor products. The examples we have in mind are spin systems, so that $A_{x}\cong M_{n}(\mathbbm{C}))$ or $\mathbbm{C}^{n}$ for some $n$. However, all the math works virtually unchanged more generally.
@@
For $F\subseteq G$, we have a natural inclusion $A_{F}\hookrightarrow A_{G}$, given by $a\mapsto 1^{\otimes G-F}\otimes a$. We define the C*-algebra
@@
While spin systems are the main example we consider, most of the definitions we provide make sense (and may be of interest) in greater generality, in particular for nets of symmetric operators in an ordinary spin system with a gauge symmetry. The following definition, which abstracts some of the main features of spin systems we described above, has been used recently in the study of symmetric quantum cellular automata and in the theory of local topological order [Jon24], [JL24, JNPW23, TWH+23]. It is essentially a discrete metric version of an algebraic quantum field theory (AQFT) in the sense of Haag and Kastler [HK64, Haa96], but without a distinguished Hilbert space representation or symmetry group.
@@
If we are thinking of a discrete net of C*-algebras as generalizing a spin system, the major difference (aside from the type of the local C*-algebra) is witnessed by algebraic entanglement. For a general discrete net $A$, for $F\cap G=\varnothing$, our inclusions gives a natural map $A_{F}\otimes A_{G}\mapsto A_{F\cup G}$. In an ordinary spin system, this is an isomorphism of algebras. For a general net, this may not be surjective. In other words, local operators may not factorize as sums of products of operators localized in $F$ and $G$ respectively. Physically, this means that observables are not determined by the observables localized at points. See Section 4.4 for further discussion. Below we give some naturally occurring examples of discrete nets beyond ordinary spin systems.
@@ 6. Locality preserving channels
Since $(\widetilde{C}_{\phi})_{H\circ E}$ has controlled decay with respect to $\mathcal{E}$, there exists some $F$ (which we can assume contains $H$ by taking union) such that for $x,y\notin F$, $(\widetilde{C}_{\phi})_{H}(x,y)<\epsilon-(10\delta+\delta^{2})$. Thus for $(x,y)\notin F$, and by thinking of $\phi$ as a locality presreving channel to the scalars (i.e thinking of the scalars as a trivial discrete net of algebras), it follows from Theorem 6.5 that
$(\widetilde{C}_{\phi\circ\Psi})_{E}(x,y)<\epsilon.$
@@ 7.1. Topological order: superselection sectors
A Hilbert space representation $H$ of $A$ satisfies the superselection criterion if for any cone $\mathcal{D}$ in $\mathbbm{Z}^{n}$, $H$ is localizble in $\mathcal{D}$. We denote by $\text{Sec}_{\phi}$ the W*-category of Hilbert space representations of $A$ satisfying the superselection criteria.
We note that for many states (in fact, all ground states of gapped Hamiltonians) we can replace $\prec$ in the above definition with unitary equivalence without loss of generality, and this recovers the usual definition. $\text{Sec}_{\phi}$ is unitarily Cauchy complete W*-category.
Our goal in this section is to define a version of $\text{Sec}_{\phi}$ which makes for a spin system over any metric space $(X,d)$ and depends only on its quasi-isometry class.
So let $(X,d)$ be a metric space with finite asymptotic dimension $n$, and suppose $(X,d)$ is quasi-geodesic. A coarse $L,C$ cone is the image of a quasi-isometric embedding from a cone $\mathcal{D}\in\mathbbm{R}^{n}$ onto a subspace of $(X,d)$, with Large-scale Lipschitz constants $L,C$. The asymptotic dimension assumption is simply to guarantee we are using the “right” dimension of cone for the region, though it is not strictly necessary below. Recall also that for a subset $E\subseteq X$, $E+R$ is simply the $R$-neighborhood of $E$ and does not refer to translation by $R$ which does make sense in general.
@@
A Hilbert space representation of $A$ satisfies the coarse superselection criterion on $(X,d)$ if for any coarse $L,C$ cone $\mathcal{D}$, there exists some $R_{L,C}\geq 0$ such that $H$ is localizable in $\mathcal{D}+R_{L,C}$. We denote by $\text{Sec}^{d}_{\phi}$ the W*-category of Hilbert space representations of $A$ satisfying the coarse superselection criteria.
@@
Then any coarse $K,B$ cone with respect to $d$ is a coarse $LK,LB+C$ cone with respect to $d^{\prime}$ and conversely, $K,B$ cones with respect to $(X,d^{\prime})$ are $L^{-1}K,L^{-1}B-C$ cones with respect to $d$. Now let $H\in\text{Sec}^{d}_{\phi}$, and pick a coarse $K,B$ cone $\mathcal{D}$ with respect to $\mathcal{D}^{\prime}$. Then there is some $R_{L^{-1}K,L^{-1}B-C}\geq 0$ such that $H$ is localizble in $\mathcal{D}+R_{L^{-1}K,L^{-1}B-C}$. Then for any choice of $R^{\prime}_{K,B}>L(R_{L^{-1}K,L^{-1}B-C})+C$, then $d^{\prime}(x,\mathcal{D})>R^{\prime}_{K,B},$ which implies $d(x,\mathcal{D})\geq L^{-1}(R^{\prime}_{K,B}-C)>R_{L^{-1}K,L^{-1}B-C}$. Thus
$(\mathcal{D}+R^{\prime}_{K,B})^{c}\subseteq(\mathcal{D}+R_{L^{-1}K,L^{-1}B-C})^{c},$
where the $+R^{\prime}_{K,B}$ is defined with resect to the $d^{\prime}$ metric and the $+R_{L^{-1}K,L^{-1}B-C}$ with respect to the $d$-metric. Therefore $A_{(\mathcal{D}+R^{\prime}_{K,B})^{c}}$ is contained, as a subalgebra, of $A_{(\mathcal{D}+R_{L^{-1}K,L^{-1}B-C})^{c}}$ and the superselction criterion restricts to
$H|_{A_{(\mathcal{D}+R^{\prime}_{K,B})^{c}}}\preceq(H_{\phi})|_{A_{(\mathcal{D}+R^{\prime}_{K,B})^{c}}}.$

### v2

@@ Abstract.
The prospect of realizing highly entangled states on quantum processors with fundamentally different hardware geometries raises the question: to what extent does a state of a quantum spin system have an intrinsic geometry? In this paper, we propose that both states and dynamics of a spin system have a canonically associated coarse geometry, in the sense of Roe, on the set of sites in the thermodynamic limit. For a state $\phi$ on an (abstract) spin system with an infinite collection of sites $X$, we define a universal coarse structure $\mathcal{E}_{\phi}$ on the set $X$ with the property that a state has decay of correlations with respect to a coarse structure $\mathcal{E}$ on $X$ if and only if $\mathcal{E}_{\phi}\subseteq\mathcal{E}$. We show that under mild assumptions, the coarsely connected completion $(\mathcal{E}_{\phi})_{con}$ is stable under quasi-local perturbations of the state $\phi$. We also develop in parallel a dynamical coarse structure for arbitrary quantum channels, and prove a similar stability result. We show that several order parameters of a state only depend on the coarse structure of an underlying spatial metric, and we establish a basic compatibility between the dynamical coarse structure associated to a quantum circuit $\alpha$ and the coarse structure of the state $\psi\circ\alpha$ where $\psi$ is any product state.
@@ 1. Introduction
Recently, there have been remarkable experimental realizations of highly entangled, topologically ordered ground states of gapped quantum spin liquids on quantum processors ([S+21, I+24, Xo24, A+23]). The states constructed in these experiments have (at least in principle) the same entanglement structure as the ground state of the theoretical spin liquid up to local perturbation. Thus, properties of the spin liquid which emerge from the entanglement structure of the ground state alone, and are robust against local perturbation, are arguably realized in the processor rather than merely simulated. A key example of such a property is topological order [Kit06, ZCZW19],
Many of the interesting properties that we think of as emerging from the entanglement structure of a state also depend on the geometry of the lattice sites to define mathematically and in practice. In most of the experiments realizing topological order that have been performed so far, the spatial geometry of the processor’s hardware roughly agrees with the geometry of the system being realized. However, for many states of interest, the geometry underlying the processor’s hardware is neccessarily incompatible with the geometry of the system from which the state originates (which occurs to some extent in the experiment [I+24]).
To illustrate the problem, suppose that we prepare the ground state of a 3D spin liquid (e.g. a Walker-Wang model [WW12]) on a quantum processor with 2D hardware, making use of some (necessarily non-local) bijective encoding scheme to identify sites of the 3D lattice with sites in the 2D processor111in principle we could imagine a more complicated encoding scheme which doesn’t directly map sites to sites, but we will see that allowing this can dramatically change the universal features of a system. If we, as the experimenter, remember this encoding scheme, then it seems that the geometric incompatibility between the two systems may not be an issue, since the universal properties depending on the 2D metric of the system are still mathematically well defined as long as we utilize our encoded 3D metric. However, if we are simply handed the state on the processor and not told the encoding scheme, is it possible in principle to work out the “correct” geometry on the set of sites to be used when defining and measuring characteristics such as topological order, independently of the physical spatial relations of the local Hilbert spaces?
@@
Is there a precise mathematical sense in which the “correct” geometry on the set of sites of a spin system, with respect to which entanglement-based order parameters should be defined, is intrinsic to a state? Or do we need to rely on the extrinsic spatial geometry of the sites to define entanglement-based order parameters?
Another theoretical motivation for the above question comes from making sense of various notions of equivalence of quantum phases. Traditionally, we say ground states of two gapped Hamiltonians are in the same quantum phase if there is a (smooth, continuous, etc.) path of Hamiltonians between the two that keeps the gap open. However, the quantum information perspective has motivated the consideration of state-based definitions of equivalence that do not require reference to a parent Hamiltonian. Examples include finite-depth circuit equivalence, locality-preserving equivalence [CGW10, ZCZW19], and quasi-local automorphic equivalence [BMNS12]. We stress that all of these notions depend not just on the state, but a-priori also on the underlying geometry of the lattice to define. Thus the“intrinsic geometry” of a state should tell you which geometric structure you should be using when trying to characterize quantum phases.
@@
For a fixed set $X$, it will be useful for us to think of the collection of all coarse structures on it as a partially ordered set with respect to containment. Now note that two metrics on $X$ induce the same coarse structure precisely when they have the same controlled set, and having the same coarse structure will be the notion that is most relevant to use, rather than abstract coarse equivalence.
@@
$(\mathcal{E}_{\phi})_{con}$ in the above Theorem can be viewed as a “connected completion” of $\mathcal{E}_{\phi}$ generated by adding all finite subsets of $X\times X$ to $\mathcal{E}_{\phi}$.
The above theorem gives a complete characterization of (connected) coarse structures with respect to which $\phi$ is local, and gives natural candidates for a choice of “universal coarse geometry”. There are other natural variations of $\mathcal{E}_{\phi}$ that are similarly universal but with additional hypotheses. The main variation we consider in this paper involves higher order controlled decay, which gives rise to the universal coarse structure $\mathcal{E}_{\widetilde{\phi}}$ (see Definition 4.7 and Theorem 4.8). In this case, instead of considering operators localized at points, we consider operators localized in uniformly bounded regions. For resonable nice states, we would expect all these variations to agree. The reason we consider all these variations here is that they make the statement of some of our theorems more convenient.
@@ 1.2. Stability under quasi-local perturbation
Let $A$ be an (abstract) spin system over the set $X$, and $\alpha:A\rightarrow A$ be any ucp map such that $\mathcal{E}_{\alpha}$ is proper and monogenic. Then for any quasi-local perturbation $\Psi:A\rightarrow A$, $(\mathcal{E}_{\alpha\circ\Psi})_{con}=(\mathcal{E}_{\alpha})_{con}$.
@@ 1.4. Coarse dependence of order parameters
In the final section of the paper, we consider the observation that many entanglement-based order parameters of a state, such as topological order, depend explicitly on some kind of geometry to be defined. We have proposed (a family of) candidates for canonical coarse geometries that need not be compatible with spatial geometries directly, but we need to work out the extent that order parameters of interest only depend on coarse geometries in the first place. This is the topic of Section 7.
@@
The goal, then, is to see which definitions of various order parameters, depending a-priori on a metric $d$, are invariant under this equiavalence relation.
Superselection sectors and topological order. The first order parameter we investigate in Section 7.1 is the W*-category of cone-localized superselection sectors of a state $\phi$ defined on a spin system in $\mathbbm{Z}^{n}$. In 2-dimensions, under some assumptions this category can be upgraded to a braided tensor category, which characterizes the states’ topological order [Naa11, Oga22]. For $(X,d)$ having finite asymptotic dimension (see Definition 4.9), we give a coarse variation of the definition of superselction sectors in terms of coarse cones, and we show that this W*-category only depends on the coarse structure. We also show that in the case that the space is $\mathbbm{Z}^{n}$ with the Euclidean metric, this agrees with the usual definition of superselction sectors in terms of localization in honest cones (see Proposition 7.4), hence this W*-category is a coarse order parameter.
@@
(c.f. Proposition 7.6). The definitions correlation length 00, finite correlation length, and infinite correlation length for a state $\phi$ each depend only on the quasi-equivalence class of the metric. Thus if $\mathcal{E}_{\phi}$ is monogenic, then having 00, finite, or infinite correlation length is a coarse order parameter.
@@ 1.7. Acknowledgements
We would like to thank Pieter Naaijkens, David Penneys and Dominic Williamson for interesting and useful discussions around these topics. Both authors were supported by NSF DMS- 2247202.
@@ 2. C*-algebra formalism
The mathematical setup for this paper uses the C*-algebra framework for describing physical systems. We briefly review this here, but refer the reader to [BR87, Lan17] for all definitions and results we mention below concerning C*-algebras, as well as a more extensive overview of their role in mathematical physics.
@@
A single automorphism $\alpha\in Aut(A)$ is interpreted as a discrete time evolution. If a system is in state $\phi$ at time 00, then after one time step, it is in state $\phi\circ\alpha$. A continuous time evolution is decribed by a homomorphism $\mathbbm{R}\rightarrow\text{Aut}(A)$, $t\mapsto\alpha_{t}$. In this case, if a system is initially in state $\phi$ at time $t=0$, then at time $t$ the state is $\phi\circ\alpha_{t}$. If the homomorphism is continuous in the point norm topology, $\{\alpha_{t}\}$ has an infinitesimal generator $\delta$ which is a (generally unbounded) derivation on the C*-algebra $A$. This can be identified as the Hamiltonian of the system, and the equation expressing that time evolution is generated by $\delta$ is a generalization of the Schrodinger equation (for further details, see [BR87]).
More generally, we may consider open dynamics, which are given instead by quantum channels. These are expressed in the Heisenberg picture as unital, completely positive maps $\alpha:A\rightarrow A$, and the action on states is $\phi\mapsto\phi\circ\alpha$ (see [Pau02]). Clearly, automorphisms (which we will call “unitary dynamics”) give examples of quantum channels. However, channels actually simultaneously generalize states and dynamics, since any state also defines a channel, via the map $a\mapsto\phi(a)1_{A}$. In this paper, we will be thinking of a unital completely positive maps mostly as generalized dynamics.
@@ 2.1. Discrete nets of C*-algebras
. The fundamental example of a quantum system where the algebraic framework can be very useful is the thermodynamic limit of a spin system [BR97]. Here, we start with a (usually countably infinite) set of sites $X$. At each site $x\in X$ we have a unital C*-algebra $A_{x}$ of localized observables. In this paper, we will assume this algebra is nuclear, so that we can ignore the technicalities arises from the different types of tensor products. The examples we have in mind are spin systems, so that $A_{x}\cong M_{n}(\mathbbm{C}))$ or $\mathbbm{C}^{n}$ for some $n$. However, all the math works virtually unchanged more generally.
@@
This means we take the tensor product algebra whose tensor factors are explicitly indexed by the elements of $F$. For $F\subseteq G$, we have a natural inclusion $A_{F}\hookrightarrow A_{G}$, given by $a\mapsto 1^{\otimes G-F}\otimes a$. We define the C*-algebra
@@
While spin systems are the main example we consider, most of the definitions we provide make sense (and may be of interest) in greater generality, in particular for nets of symmetric operators in an ordinary spin system with a global gauge symmetry. The following definition, which abstracts some of the main features of spin systems we described above, has been used recently in the study of symmetric quantum cellular automata and in the theory of local topological order [Jon24], [JL24, JNPW23, TWH+23]. It is essentially a discrete metric version of an algebraic quantum field theory (AQFT) in the sense of Haag and Kastler [HK64, Haa96], but without a distinguished Hilbert space representation or symmetry group.
@@
If we are thinking of a discrete net of C*-algebras as generalizing a spin system, the major difference (aside from the type of the local C*-algebra) is witnessed by algebraic entanglement. For a general discrete net $A$, for $F\cap G=\varnothing$, our inclusions gives a natural map $A_{F}\otimes A_{G}\mapsto A_{F\cup G}$. In an ordinary spin system, this is an isomorphism of algebras. For a general net, this may fail to be injective or surjective. In other words, products of operators localized in disjoint regions may be 00, and local operators may not factorize as sums of products of operators localized in $F$ and $G$ respectively. Physically, this latter condition implies that local observables are not completely determined by the observables localized at points. See Section 4.4 for further discussion. Below we give some naturally occurring examples of discrete nets beyond ordinary spin systems.
@@ 6. Locality preserving channels
and the same bound for $a^{\prime}$ and $b^{\prime}$. Since $(\widetilde{C}_{\phi})_{H\circ E}$ has controlled decay with respect to $\mathcal{E}$, there exists some $F$ (which we can assume contains $H$ by taking union) such that for $x,y\notin F$, $(\widetilde{C}_{\phi})_{H}(x,y)<\epsilon-(10\delta+\delta^{2})$. Thus for $(x,y)\notin F$, and by thinking of $\phi$ as a locality presreving channel to the scalars (i.e thinking of the scalars as a trivial discrete net of algebras), it follows from Theorem 6.5 that
$(\widetilde{C}_{\phi\circ\Psi})_{F}(x,y)<\epsilon.$
@@ 7.1. Topological order: superselection sectors
A Hilbert space representation $H$ of $A$ satisfies the superselection criterion if for any cone $\mathcal{D}$ in $\mathbbm{Z}^{n}$, $H$ is localizable in $\mathcal{D}$. We denote by $\text{Sec}_{\phi}$ the W*-category of Hilbert space representations of $A$ satisfying the superselection criteria.
We note that for many states (in fact, all ground states of gapped Hamiltonians) we can replace $\prec$ in the above definition with unitary equivalence without loss of generality, and this recovers the usual definition. $\text{Sec}_{\phi}$ is a unitarily Cauchy complete W*-category [GLR85].
Our goal in this section is to define a version of $\text{Sec}_{\phi}$ which makes sense for a spin system over any metric space $(X,d)$ and depends only on its quasi-isometry class.
So let $(X,d)$ be a metric space with finite asymptotic dimension $n$, and suppose $(X,d)$ is quasi-geodesic. Let $\mathcal{D}$ be a cone in $\mathbbm{R}^{n}$. For $L>0,C\geq 0$, a coarse $L,C,\mathcal{D}$ cone is a quasi-isometric embedding $f:\mathcal{D}\rightarrow(X,d)$, with Large-scale Lipschitz constants $L,C$. The asymptotic dimension assumption is simply to guarantee we are using the “right” dimension of cone for the region, though it is not strictly necessary below. Recall also that for a subset $E\subseteq X$, $E+R$ is simply the $R$-neighborhood of $E$ and does not refer to translation by $R$ which does make sense in general.
@@
A Hilbert space representation of $A$ satisfies the coarse superselection criterion on $(X,d)$ if for any triple $L,C,\mathcal{D}$, there exists some constant $R\geq 0$ such that for an coarse $L,C,\mathcal{D}$ cone $f:\mathcal{D}\rightarrow X$, $H$ is localizable in $f(\mathcal{D})+R$. We denote by $\text{Sec}^{d}_{\phi}$ the W*-category of Hilbert space representations of $A$ satisfying the coarse superselection criteria.
@@
Then any coarse $K,B,\mathcal{D}$ cone with respect to $d$ is a coarse $LK,LB+C,\mathcal{D}$ cone with respect to $d^{\prime}$ and conversely, $K,B,\mathcal{D}$ cones with respect to $(X,d^{\prime})$ are $L^{-1}K,L^{-1}B-C,\mathcal{D}$ cones with respect to $d$. Now let $H\in\text{Sec}^{d}_{\phi}$, and let $K,B,\mathcal{D}$ and $f:\mathcal{D}\rightarrow(X,d^{\prime})$ be a coarse cone. Then there is some $R\geq 0$ such that $H$ is localizble in $f(\mathcal{D})+R$. Then for any choice of $R^{\prime}>LR+C$,
$d^{\prime}(x,f(\mathcal{D}))>R^{\prime}$ which implies $d(x,f(\mathcal{D}))\geq L^{-1}R^{\prime}-C>R$. Thus
$(f(\mathcal{D})+_{d^{\prime}}R^{\prime})^{c}\subseteq(f(\mathcal{D})+_{d}R)^{c},$
Therefore $A_{(\mathcal{D}+_{d^{\prime}}R^{\prime})^{c}}$ is contained in $A_{(\mathcal{D}+_{d}R)^{c}}$ and the superselction criterion restricts to
$H|_{A_{(\mathcal{D}+_{d^{\prime}}R^{\prime})^{c}}}\preceq(H_{\phi})|_{A_{(\mathcal{D}+_{d^{\prime}}R^{\prime})^{c}}}.$